# A Survey of Multimodal Large Language Model from A Data-centric Perspective

### v1

@@ 2.3. Data-Centric AI and Why Data-Centric LLMs
Analyzing LLMs and MLLMs from a data-centric perspective offers several advantages. First, it enables researchers to identify and address potential biases and limitations in the training data that can propagate to the models’ outputs. By carefully curating and augmenting datasets to ensure diversity and representativeness, researchers can mitigate biases and improve the fairness and generalizability of the models (Chen et al., 2024a). Second, a data-centric approach allows for a deeper understanding of the models’ capabilities and limitations based on the characteristics of the training data (McKinzie et al., 2024). By systematically varying data properties and evaluating model performance, researchers can gain insights into the specific data attributes that contribute to the models’ success or failure in various tasks. This understanding can guide the development of more efficient and effective data collection and curation strategies.
Moreover, adopting a data-centric perspective in the development of LLMs and MLLMs opens up opportunities for data-efficient learning and adaptation (Liu et al., 2024b, 2023c; Chen et al., 2023e). Techniques such as few-shot learning and prompt engineering enable models to adapt to new tasks or domains with limited task-specific data. This adaptability is particularly valuable in scenarios where labeled data is scarce or expensive to obtain. Additionally, data-centric approaches facilitate the development of more interpretable and explainable models by linking model behaviors to specific data characteristics, allowing for the identification of relevant data subsets for debugging and error analysis.
The shift towards data-centric AI has underscored the crucial role of data in the development and evaluation of LLMs and MLLMs. By prioritizing data quality, diversity, and representativeness, researchers can build more robust, unbiased, and adaptable models. Analyzing these models from a data-centric perspective offers valuable insight into their capabilities, limitations, and potential for improvement. As the field continues to advance, a data-centric approach will be essential to develop LLMs and MLLMs that can effectively understand, generate, and reason across multiple modalities while ensuring fairness, interpretability, and generalizability.
@@ 2.4. Categories of Data-Centric MLLMs
In this survey, we review previous work on MLLMs from a data-centric perspective. Our taxonomy is primarily based on the different stages of MLLM development, as illustrated in Figure 2. Specifically, we organize our article around a data pipeline, shown in Figure 1, where the data pipeline for MLLMs comprises three stages: pre-processing (including data collection and data processing in Section 3), pre-training data processing (Section 4), and adaptation data processing (Section 5).
Along this pipeline, we discuss data-related model performance and analyze how dataset curation affects model outcomes. During the pre-processing stage, large-scale multimodal data are collected from various sources and constructed into datasets. Processing steps such as filtering, deduplication, and data enhancement are applied to improve data quality. In the pre-training stage, data selection is performed to choose appropriate data for model training. In the adaptation stage, we focus on data processing and generation for supervised fine-tuning and human preference alignment. Furthermore, we discuss data evaluation metrics and evaluation datasets for MLLMs, examining their construction and main characteristics. Finally, we suggest possible future directions for further research on data-centric MLLMs.
@@ 3. Data collecting and processing
The first and fundamental step in training MLLMs is to collect and process sufficient data from various sources. In this section, we first introduce the sources for data collection. Next, we summarize the processing steps for multimodal data, including commonly used filtering, deduplication, and data enhancement methods. Finally, we provide an overview of commonly used datasets for MLLMs. These datasets are utilized in the pre-training, adaptation, and evaluation stages of MLLMs.
@@ 3.1. Data Collecting Sources
We first introduce the sources where researchers typically find raw data for MLLM pre-training. Commonly used data sources can be divided into six categories, as shown in Figure 3: common webpages, social media, academic papers, books, code repositories, and professional sources. Data from these different sources have various characteristics, each contributing uniquely to improving different abilities of the models.
Figure 3. Commonly used data sources.
[…]
[Commonly used data sources.]Commonly used data sources.
@@
Apart from CommonCrawl, several works focus on selecting specific webpages and crawling data independently. For textual datasets, initiatives like WuDaoCorpora (Yuan et al., 2021) crawl thousands of web pages to construct vast textual corpora. In the multimodal dataset sector, there are three main self-crawling approaches: general, focused on specific platforms, and querying search engines for images. Examples include AI Challenger Captions (Wu et al., 2017) and Wukong (Gu et al., 2022). General approaches, such as CC3M (Sharma et al., 2018) and CC12M (Changpinyo et al., 2021), involve crawling billions of web pages without targeting specific platforms. Focused approaches, on the other hand, involve selecting specific webpages for crawling. For example, Wikipedia is a commonly used resource that serves as a rich and accurate source for text-based single-modal pre-training datasets like the Pile (Gao et al., 2020). It also plays a crucial role in creating multimodal pre-training datasets such as WIT (Srinivasan et al., 2021) and WikiCaps (Schamoni et al., 2018).
For the vision modality, Flickr (Young et al., 2014) serves as an important source, containing photos and videos shared by online users. This contributes to image-text datasets such as Flickr30k (Young et al., 2014). For the audio modality, platforms like BBC Sound Effects, FreeSound, and SoundBible are commonly used, contributing to datasets such as WavCaps (Mei et al., 2023) and FSD50K (Fonseca et al., 2022).
@@
Social media enriches MLLM training datasets with real-time, varied, and colloquial text data, as well as multimodal content, capturing human expression and interaction. Stack Exchange is a vital source of high-quality textual and visual content, hosting extensive forums with rich dialogues and images. This contributes to text-only datasets such as SlimPajama (Soboleva et al., 2023) and image-text datasets such as VQAonline (Chen et al., 2023g). Beyond Stack Exchange, Reddit features user-shared content with discussions and voting on a wide array of topics, contributing to text-rich datasets such as Pushshift Reddit (Baumgartner et al., 2020) and visual datasets such as RedCaps (Desai et al., 2021) and MREDDITSUM (Overbay et al., 2023).
For video data, YouTube stands out as a video-sharing platform with over 500 hours of video uploaded per minute. It offers a vast dataset of multimodal information, contributing to datasets such as HowTo100M (Miech et al., 2019), which leverages instructional videos for multimodal learning, and YouCook2 (Zhou et al., 2017), focused on cooking videos. X, commonly known by its former name, Twitter, is a popular social media platform for sharing text messages, images, audio, and videos. Datasets derived from X include TweetEval (Barbieri et al., 2020), designed for evaluating sentiment and stance.
@@
Academic papers offer high-quality, authoritative content, enriching pre-training datasets with specialized knowledge and formal language, which is critical for developing language models with expertise in academic and professional domains. Currently, academic papers are primarily utilized as sources for text-based pre-training datasets.
arXiv serves as a crucial data source, offering a vast array of prepublished scientific research for training specialized language models. This is evidenced by its role in creating datasets like RedPajama-Data-1T (Computer, 2023), Pile-arXiv (Gao et al., 2020), and image-text datasets such as MMC (Liu et al., 2023d).
S2ORC (Lo et al., 2020) is a corpus containing 81.1 million academic papers, offering metadata, abstracts, references, and full texts for 8.1 million open-access works. It compiles materials from various sources, forming the largest publicly accessible academic dataset, such as peS2o (Soldaini and Lo, 2023) and Dolma (Soldaini et al., 2024).
@@
Project Gutenberg acts as a pivotal source for language modeling and analysis with its extensive library of more than 70,000 free eBooks, fueling datasets such as PG-19 (Rae et al., 2019). The BookCorpus (Zhu et al., 2015), sourced from Smashwords, encompasses a diverse collection of 11,038 self-published novels that span genres such as romance, science fiction, and fantasy. This extensive corpus has been instrumental in training landmark models such as GPT (Radford et al., [n. d.]) and BERT (Kenton and Toutanova, 2019).
Bibliotik consists of a mix of fiction and non-fiction books. Books3 (Gao et al., 2020), part of the Pile dataset, is derived from the Bibliotik private tracker made available by Shawn Presser in 2020, containing about 197,000 books. This dataset is essential for context modeling and narrative research, accounting for 2.1% of the RedPajama-Data-1T dataset.
@@
Code.
Training LLMs on extensive code datasets significantly boosts their capability to generate high-quality, accurate programs, as evidenced by their enhanced performance in expert-designed tests and complex programming challenges. The most common sources for codes are BigQuery and GitHub. BigQuery, a Google data warehouse, contributes to code datasets such as BIGQUERY, which encompasses open-source code in six programming languages: C, C++, Go, Java, JavaScript, and Python. This dataset is aimed at training multilingual CODEGEN models (CODEGEN-MULTI) (Nijkamp et al., 2022). GitHub, a premier platform for software collaboration, supports datasets like The Stack (Kocetkov et al., 2022), BigPython (Nijkamp et al., 2022), and the Pile (Gao et al., 2020) with its vast, open-source repositories and global developer community. The Stack (Kocetkov et al., 2022) leverages GitHub’s permissively licensed code and governance tools to support the StarCoder (Li et al., 2023a) project with a diverse programming language dataset.
@@
To enhance the performance of LLMs in specific domains, general pre-training is supplemented with incremental, domain-specific pre-training. In the legal domain, pre-training data is primarily textual. FreeLaw stands out by offering free access to judgments from India’s higher courts, exemplified by the Pile-FreeLaw (Gao et al., 2020) dataset, which structures these judgments for machine learning.
In the math domain, textual datasets primarily originate from Khan Academy’s exercises for foundational concepts and DeepMind’s Pile-DeepMind Mathematics (Saxton et al., 2018) for advanced, algorithmically generated problems. Additionally, DVQA (Kafle et al., 2018) introduces a multimodal aspect by combining bar charts with question-answer pairs generated through Matplotlib.
In the medical domain, pre-training data comes from online medical websites, knowledge bases, and in-hospital database systems. Online medical websites such as Qianwen Health and PubMed contribute to datasets like Huatuo-26M (Li et al., 2023j) and MedQuAD (Ben Abacha and Demner-Fushman, 2019). Knowledge bases like Wikipedia also contribute to Huatuo-26M (Li et al., 2023j) and MedHop (Welbl et al., 2018). In-hospital database systems, including electronic health record systems (EHR) and ICU-specific clinical information systems, are used in MIMIC-IV (Johnson et al., 2023). Multimodal datasets in radiology, such as MIMIC-CXR-JPG (Johnson et al., 2019) and PADCHEST (Bustos et al., 2020), are generated from in-hospital radiology reports.
In the financial domain, pre-training primarily relies on textual data sourced from both English and Chinese platforms. English contributions come from EDGAR and the SEC Financial Statement and Notes Data Sets, which enhance the FinTree project (Ok, 2023). Chinese sources include major sites like Sina Finance and Tencent Finance, as well as Eastmoney for specific documents, and forums like Guba and Xueqiu, culminating in the comprehensive BBT-FinCorpus (Lu et al., 2023b). This assortment of sources underscores the rich textual foundation for financial model pre-training.
@@ 3.2.1. Filtering
Data filtering in MLLM training is critical for enhancing model reliability and efficiency. Filtering out undesirable content is essential, as even minimal exposure to hate speech can negatively influence model behavior (Luccioni and Viviano, 2021; Gunasekar et al., 2023). For multimodal datasets, previous work has focused on filtering out unwanted data from different modalities separately. Specifically, for X-Text datasets, they consider filtering data based on both the text and the X modality independently.
Textual filters primarily encompass language and content filtration. In language filtering, documents or sentences that fall below a certain threshold for a specific language are removed. For English-only datasets, commonly used tools like Langdetect and FastText operate at the document level. For instance, C4 (Raffel et al., 2020) utilizes Langdetect to filter out any non-English pages with a probability below 0.99, reducing the Common Crawl April 2019 corpus from 6TB to 20GB (c4/en.noclean). Another renowned English language filter, FastText, employed by Dolma (Soldaini et al., 2024), eliminates documents classified as English with a probability less than 0.5, filtering out over 60% of documents. RefinedWeb (Penedo et al., 2023) utilizes a higher FastText threshold (0.65) to obtain purer English documents. For Chinese datasets, WuDaoCorpora (Yuan et al., 2021) implements heuristic rules to filter out web pages containing fragments with more than ten successive non-Chinese characters at the document level. CLUECorpus2020 (Xu et al., 2020) adopts a sentence-level filtering approach, selecting sentences whose language type is Chinese if a language is mentioned. For multilingual datasets, FastText, trained on Wikipedia data to classify 176 languages, remains a preferred choice, capable of processing 1,000 documents per second on a single CPU core. ROOTS (Laurençon et al., 2022) utilizes FastText for document-level language classification, resulting in a 1.6TB dataset containing 59 languages. For code datasets, detection methods are more elementary, often based on file extensions. For instance, in Palm (Chowdhery et al., 2023), files are filtered based on filename extensions to restrict to one of 24 common programming languages, resulting in 196GB of source code.
Another type of textual filtering is content filtering. Content filtering includes removing toxic and distracting content. Toxic content filtering targets text that is deemed rude, disrespectful, or unreasonable, employing heuristic and rule-based methods. For instance, in C4 (Raffel et al., 2020), pages containing words listed as ”Dirty, Naughty, Obscene, or Otherwise Bad Words” are removed. Distracting content includes short or incomplete sentences, ”lorem ipsum” text, and useless identifiers like HTML, CSS, and JavaScript tags. Filtering methods may involve multiple rule-based filters, such as punctuation-based segmentation in WudaoCorpora (Yuan et al., 2021), or machine learning approaches, as seen in phi-1 (Gunasekar et al., 2023). In phi-1, GPT-4 annotates a small portion of high-quality data, which is then used to train a random forest classifier for high-quality content classification.
@@
For video-level filtering, previous work has leveraged image or image-text filtering methods to handle static content and applied specific techniques for dynamic content. Video filtering involves four main components: scene transition detection, video quality and integrity improvement, refinement and coherence evaluation, and modality completeness. For scene transition detection, algorithms are used to remove abrupt scene changes, ensuring a cleaner dataset free from disruptive transitions (Wang et al., 2023a; Blattmann et al., 2023). To improve video quality and integrity, three aspects are considered: motion, textual information in the video, and resolution and frame rate. For motion analysis, dense optical flow (Farnebäck, 2003) is typically utilized to measure motion complexity and filter out static or repetitive sequences with limited educational value. To detect textual information in videos, OCR (Baek et al., 2019) is applied to identify and remove clips with substantial text, which can otherwise misguide the model’s visual interpretation. The resolution and frame rate of training video datasets are standardized, enabling VLMs to learn from uniformly structured data (Wang et al., 2023a). In constructing video-text datasets, ensuring the relevance and clarity of annotations is paramount. Focused efforts are made to refine the textual content of video-text pairs, particularly subtitles, to ensure they are contextually meaningful and devoid of promotional or irrelevant material (Xu et al., 2023c). Models such as CLIP are employed to evaluate the coherence between video frames and the accompanying text (Chen et al., 2024b; Xu et al., 2023c; Wang et al., 2023a), ensuring that annotations accurately reflect the video content. In some studies, clips missing any modalities—vision, audio, or subtitles—are excluded to meet the requirements for comprehensive understanding across all modalities (Chen et al., 2024b).

### v2

@@ 2.3. Data-Centric AI and Why Data-Centric MLLMs
Analyzing LLMs and MLLMs from a data-centric perspective offers several advantages. First, it enables researchers to identify and address potential biases and limitations in the training data that can propagate to the models’ outputs. By carefully curating and augmenting datasets to ensure diversity and representativeness, researchers can mitigate biases and improve the fairness and generalizability of the models (Chen et al., 2024a). Second, a data-centric approach allows for a deeper understanding of the models’ capabilities and limitations based on the characteristics of the training data (McKinzie et al., 2024). By systematically varying data properties and evaluating model performance, researchers can gain insights into the specific data attributes that contribute to the models’ success or failure in various tasks. This understanding can guide the development of more efficient and effective data collection and curation strategies. Moreover, adopting a data-centric perspective in the development of LLMs and MLLMs opens up opportunities for data-efficient training and adaptation (Liu et al., 2024b, 2023c; Chen et al., 2023e). Additionally, data-centric approaches facilitate the development of more interpretable and explainable models by linking model behaviors to specific data characteristics.
@@ 2.4. Categories of Data-Centric MLLMs
In this survey, we review previous work on MLLMs from a data-centric perspective. Our taxonomy is primarily based on different stages of MLLM development, as illustrated in Figure 2. Specifically, we organize our article around a data pipeline, shown in Figure 1, where the data pipeline for MLLMs comprises three stages: pre-processing (including data collection and data processing in Section 3), pre-training data processing in Section 4), and adaptation data processing in Section 5. Along this pipeline, we discuss data-related model performance and analyze how dataset curation affects model performance. During the pre-processing stage, large-scale multimodal data are collected from various sources and constructed into datasets. Processing steps such as filtering, deduplication, and data enhancement are applied to improve data quality. In the pre-training stage, data selection, domain mixing and modality mixing are performed to choose appropriate data for model training. In the adaptation stage, we focus on data processing and generation for supervised fine-tuning and human preference alignment. Furthermore, we discuss data evaluation metrics and evaluation datasets for MLLMs in Section 6, examining their construction and main characteristics. Finally, we suggest possible future directions for further research on data-centric MLLMs in Section 7.
@@ 3. Data collecting and processing
The first and fundamental step in training MLLMs is to collect and process sufficient data from various sources. In this section, we first introduce the sources for data collection. Next, we summarize the processing steps for multimodal data, including commonly used filtering, deduplication, and data enhancement methods. Finally, we provide an overview of commonly used datasets for MLLMs.
@@ 3.1. Data Collecting Sources
We first introduce the sources where researchers typically find raw data for MLLM pre-training. Commonly used data sources can be divided into six categories: common webpages, social media, academic papers, books, code repositories, and professional sources. Data from these different sources have various characteristics, each contributing uniquely to improving different abilities of the models.
@@
Apart from CommonCrawl, several works focus on selecting specific webpages and crawling data independently. For textual datasets, initiatives like WuDaoCorpora (Yuan et al., 2021) crawl thousands of web pages to construct vast textual corpora. In the multimodal dataset sector, there are three main self-crawling approaches: general approach, focused approach on specific platforms, and querying search engines for images. Examples include AI Challenger Captions (Wu et al., 2017) and Wukong (Gu et al., 2022). General approaches, such as CC3M (Sharma et al., 2018) and CC12M (Changpinyo et al., 2021), involve crawling billions of web pages without targeting specific platforms. Focused approaches, on the other hand, involve selecting specific webpages to crawl. For example, Wikipedia is a commonly used resource that serves as a rich and accurate source for text-based single-modal pre-training datasets like the Pile (Gao et al., 2020). It also plays a crucial role in creating multimodal pre-training datasets such as WIT (Srinivasan et al., 2021) and WikiCaps (Schamoni et al., 2018). For the vision modality, Flickr (Young et al., 2014) serves as an important source, containing photos and videos shared by online users. This contributes to image-text datasets such as Flickr30k (Young et al., 2014). For the audio modality, platforms such as BBC Sound Effects, FreeSound, and SoundBible are commonly used, contributing to datasets such as WavCaps (Mei et al., 2023) and FSD50K (Fonseca et al., 2022).
@@
Social media platforms significantly enhance the training datasets for MLLMs by providing real-time, varied, and colloquial text data, as well as multimodal content that captures human expression and interaction. However, questions regarding the ownership and copyright of social media content are on the rise. MLLM researchers need to be aware of these issues and understand the rights and licensing of the data sources they use. Stack Exchange offers high-quality textual and visual content, hosting extensive forums with rich dialogues and images. In addition to Stack Exchange, Reddit offers user-generated content with discussions and voting on a wide array of topics, contributing to both textual and visual content. For video data, YouTube is a standout platform, with a vast amount of videos uploaded every minute. X, formerly known as Twitter, offers great amount of sharing text messages, images, audio, and videos.
@@
Academic papers offer high-quality, authoritative content, enriching pre-training datasets with specialized knowledge and formal language, which is critical for developing language models with expertise in academic and professional domains. Academic papers are primarily utilized as sources for text-based pre-training datasets. arXiv offers a vast array of prepublished scientific research for training specialized language models. This is evidenced by its role in creating datasets such as RedPajama-Data-1T (Computer, 2023), Pile-arXiv (Gao et al., 2020), and image-text datasets such as MMC (Liu et al., 2023d). Another academic platform semantic scholars contribute to the S2ORC (Lo et al., 2020) corpus, which contains 81.1 million academic papers, offering metadata, abstracts, references, and full texts for 8.1 million open-access works.
@@
Project Gutenberg acts as a pivotal source for language modeling and analysis with its extensive library of more than 70,000 free eBooks, fueling datasets such as PG-19 (Rae et al., 2019). Smashwords contributes to the well-known BookCorpus (Zhu et al., 2015) dataset, which encompasses a diverse collection of 11,038 self-published novels that span genres such as romance, science fiction, and fantasy. This extensive corpus has been instrumental in training landmark models such as GPT (Radford et al., [n. d.]) and BERT (Kenton and Toutanova, 2019). Bibliotik that consists of a mix of fiction and non-fiction books contributes to Books3 (Gao et al., 2020), part of the Pile dataset, which contains about 197,000 books. This dataset is essential for context modeling and narrative research, accounting for 2.1% of the RedPajama-Data-1T dataset.
@@
To enhance the performance of MLLMs in specific domains, general pre-training is supplemented with incremental, domain-specific pre-training. In the legal domain, pre-training data is primarily textual. FreeLaw stands out by offering free access to judgments from India’s higher courts, exemplified by the Pile-FreeLaw (Gao et al., 2020) dataset, which structures these judgments for machine learning. In the math domain, textual datasets primarily originate from Khan Academy’s exercises for foundational concepts and DeepMind’s Pile-DeepMind Mathematics (Saxton et al., 2018) for advanced, algorithmically generated problems. Additionally, DVQA (Kafle et al., 2018) introduces a multimodal aspect by combining bar charts with question-answer pairs generated through Matplotlib. In the medical domain, pre-training data comes from online medical websites, knowledge bases, and in-hospital database systems. Online medical websites such as Qianwen Health and PubMed contribute to datasets like Huatuo-26M (Li et al., 2023i) and MedQuAD (Ben Abacha and Demner-Fushman, 2019). Knowledge bases like Wikipedia also contribute to Huatuo-26M (Li et al., 2023i) and MedHop (Welbl et al., 2018). In-hospital database systems, including electronic health record systems (EHR) and ICU-specific clinical information systems, are used in MIMIC-IV (Johnson et al., 2023). Multimodal datasets in radiology, such as MIMIC-CXR-JPG (Johnson et al., 2019) and PADCHEST (Bustos et al., 2020), are generated from in-hospital radiology reports. In the financial domain, pre-training primarily relies on textual data sourced from both English and Chinese platforms. English contributions come from EDGAR and the SEC Financial Statement and Notes Data Sets, which enhance the FinTree project (Ok, 2023). Chinese sources include major sites like Sina Finance and Tencent Finance, as well as Eastmoney for specific documents, and forums like Guba and Xueqiu, culminating in the comprehensive BBT-FinCorpus (Lu et al., 2023b). This assortment of sources underscores the rich textual foundation for financial model pre-training.
@@ 3.2.1. Filtering
Data filtering in MLLM training is critical for enhancing model reliability and efficiency. Filtering out unwanted content is essential, as even minimal exposure to hate speech can negatively influence model behavior (Luccioni and Viviano, 2021; Gunasekar et al., 2023). For multimodal datasets, previous work has focused on filtering out unwanted data from different modalities separately. Specifically, for X-Text datasets, they consider filtering data based on both the text and the X modality independently.
Textual filters primarily encompass language and content filtering. For language filtering, documents or sentences that fall below a certain threshold for a specific language are removed. For English-only datasets, commonly used tools like Langdetect and FastText operate at the document level. For instance, C4 (Raffel et al., 2020) utilizes Langdetect to filter out any non-English pages with a probability below 0.99. Another renowned English language filter, FastText, is widely employed by datasets such as Dolma (Soldaini et al., 2024) and RefinedWeb (Penedo et al., 2023). CLUECorpus2020 (Xu et al., 2020) adopts a sentence-level filtering approach, selecting sentences whose language type is Chinese if a language is mentioned. For multilingual datasets, FastText, trained on Wikipedia data to classify 176 languages, remains a preferred choice, capable of processing 1,000 documents per second on a single CPU core. ROOTS (Laurençon et al., 2022) utilizes FastText for document-level language classification, resulting in a 1.6TB dataset containing 59 languages. For code datasets, detection methods are more elementary, often based on file extensions. For instance, in Palm (Chowdhery et al., 2023), files are filtered based on filename extensions to restrict to one of 24 common programming languages, resulting in 196GB of source code.
Another type of textual filtering is content filtering, which includes removing toxic and distracting content. Toxic content filtering targets text that is deemed rude, disrespectful, or unreasonable, employing heuristic and rule-based methods. For instance, in C4 (Raffel et al., 2020), pages containing words in a predefined ”Dirty, Naughty, Obscene, or Otherwise Bad Words” list are removed. Distracting content includes short or incomplete sentences, ”lorem ipsum” text, and useless identifiers like HTML, CSS, and JavaScript tags. Filtering methods usually involve multiple rule-based filters, such as punctuation-based segmentation in WudaoCorpora (Yuan et al., 2021), or machine learning approaches, as seen in phi-1 (Gunasekar et al., 2023), which use GPT-4 as an annotators and then train a high-quality content classifier based on the annotated results.
@@
For video-level filtering, previous work has leveraged image or image-text filtering methods to handle static content and applied specific techniques for dynamic content. Video filtering involves four main components: scene transition detection, video quality and integrity improvement, refinement and coherence evaluation, and modality completeness. For scene transition detection, algorithms are used to remove abrupt scene changes, ensuring a cleaner dataset free from disruptive transitions (Wang et al., 2023a; Blattmann et al., 2023). To improve video quality and integrity, three aspects are considered: motion, textual information in the video, and resolution and frame rate. For motion analysis, dense optical flow (Farnebäck, 2003) is typically utilized to measure motion complexity and filter out static or repetitive sequences with limited educational value. To detect textual information in videos, OCR (Baek et al., 2019) is applied to identify and remove clips with substantial text, which can otherwise mislead the model’s visual interpretation. The resolution and frame rate of training video datasets are standardized, enabling model to learn from uniformly structured data (Wang et al., 2023a). In constructing video-text datasets, ensuring the relevance and clarity of annotations is paramount. Focused efforts are made to refine the textual content of video-text pairs, particularly subtitles, to ensure they are contextually meaningful and devoid of promotional or irrelevant material (Xu et al., 2023c). Models such as CLIP are employed to evaluate the coherence between video frames and the accompanying text (Chen et al., 2024b; Xu et al., 2023c; Wang et al., 2023a), ensuring that annotations accurately reflect the video content. In some studies, clips missing any modalities—vision, audio, or subtitles—are excluded to meet the requirements for comprehensive understanding across all modalities (Chen et al., 2024b).